OpenStudy (adi3):

Pls Help will medal and fan Graph 3sin(2(x-pi/4)+1

OpenStudy (adi3):

@Miracrown @hartnn

OpenStudy (adi3):

@rebeccaxhawaii

OpenStudy (poopsiedoodle):

is it supposed to be 3sin(2(x-pi/4)+1) or 3sin(2(x-pi/4))+1

OpenStudy (adi3):

sorry too much lag

OpenStudy (adi3):

it is the second one

OpenStudy (adi3):

@imqwerty

OpenStudy (er.mohd.amir):

3sin(2(x-pi/4)+1 =3 sin(2x-pi/2)+1 =-3 sin (pi/2-2x)+1 =-3 cos 2x+1 =-3(1-2sin ^2 x)+1 =-3+6sin ^2 x+1 =6 sin ^2x -2

OpenStudy (faiqraees):

@Er.Mohd.AMIR We have to graph not expand. So it will be a lot better if it's not expanded into a squared form.

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (mathmale):

I'd recommend that you graph y=3sin x first. Next, y=3sin(2x). Next, y=3sin[2x -pi/2). Finally, shift the whole graph upward by 1 unit. Doing this graphing step by step would most likely lead to a much clearer understanding on your part.